OpenStudy (anonymous):

Three lockers are numbered consecutively (n,n+1,n+2) in such a way that the sum of the first and last lockers is the same as twice the middle locker number. What are the locker numbers?

OpenStudy (anonymous):

n + n + 2 = 2(n + 1) 2n + 2 = 2n + 2 essentially... any three consecutive numbers works. try, (5,6,7) 5 + 7 = 2 * 6 (20,21,22) 20 + 22 = 2(21)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ty